# Network-Realized Model Predictive Control Part II: Distributed Constraint Management

### v1

@@ 5.1 Placing Constraints on Exogenous Disturbance
$\left\{\scriptsize\begin{array}[]{ll}\mathcal{D}_{d}\subseteq\mathbb{R}^{n_{d}}\,\text{s.t. }d[k]\in\mathcal{D}_{d},&\mathcal{D}_{\beta_{f}}\subseteq\mathbb{R}^{n_{u}}\,\text{s.t. }\beta_{f}[k]\in\mathcal{D}_{\beta_{f}},\\ \mathcal{D}_{\beta_{s1}}\subseteq\mathbb{R}^{n_{x}}\,\text{s.t. }\beta_{s1}[k]\in\mathcal{D}_{\beta_{s1}},&\mathcal{D}_{\beta_{s2}}\subseteq\mathbb{R}^{n_{x}}\,\text{s.t. }\beta_{s2}[k]\in\mathcal{D}_{\beta_{s2}}.\end{array}\right.$ (27)
@@
$\left\{\begin{split}\Psi_{xit}\subseteq\mathbb{R}^{n_{xi}}\text{ s.t. }\psi_{xi}[k_{0}+t]\in\Psi_{xit},\,\forall\,d_{s}[k]\in\mathcal{D}_{s},\\ \Psi_{uit}\subseteq\mathbb{R}^{n_{ui}}\text{ s.t. }\psi_{ui}[k_{0}+t]\in\Psi_{uit},\,\forall\,d_{s}[k]\in\mathcal{D}_{s}.\end{split}\right.\vspace{-3mm}$ (28)
Moving onward, we address the variables associated with the first layer subcontrollers and their implementations.
@@ 5.2 Placing Constraints on First Layer Commands
$u_{i}[k]=u_{fi}[k]+u_{s2i}[k]+S^{\top}_{{ui}}\beta_{s2}[k],\,\forall\,i\in\{1:N\}.\vspace{-3mm}$ (29)
Given the state dynamics obtained in Section 4, the problem of constraining the command signals sent to the network reduces to judiciously constraining the control effort of both the first and the second layer. Thus, for each area, we choose a pair of sets
@@
${\mathcal{U}}_{fi}:=\mathcal{U}_{i}\ominus(\,\mathcal{U}_{s2i}\oplus(S^{\top}_{{ui}}\mathcal{D}_{\beta_{s2}})),\,\forall\,i\in\{1:N\}.\vspace{-3mm}$ (31)
@@
The sets given in (30a)-(30b) are the main design parameters of the second layer and they can be assimilated with the extent to which the second layer is permitted to interfere in the dynamics of the first layer’s closed-loop system, in order to ensure that (26) is satisfied. The manner in which these sets may be chosen and its associated trade-off will be discussed in the sequel.
5.3 Placing Constraints on First Layer States
@@
We begin by inspecting the distinctive structure of the matrices which make up the realizations from (8). The latter enable us to naturally partition each of the state vectors belonging to the first layer subcontrollers into $w_{i}[k]=\tiny\begin{bmatrix}w_{r(\alpha_{ui}+1)}^{\top}[k]&\dots&w_{r(\alpha_{ui}+n_{ui})}^{\top}[k]\end{bmatrix}$, in which the vectors $w_{r\ell}[k]$ have $n_{r\ell}$ rows for all $\ell\in\{1:n_{u}\}$, i.e., the same as the matrices $A_{r\ell}$ and $B_{r\ell}$ from (10). Recalling (10), the key observation which underpins the entire mathematical mechanism presented in this section is the fact that $\mathrm{row}_{\ell}(u_{f}[k])=C_{r\ell}w_{r\ell}[k]=\mathrm{row}_{1}(w_{r\ell}[k])$, for all $\ell\in\{1:n_{u}\}$. Thus, our current objective becomes apparent: we aim to employ the bounds placed upon the partitions of $u_{f}[k]$ in order to satisfy further set inclusions with respect to $w_{r\ell}[k]$, in a distributed manner.
@@
$\mathrm{row}_{\ell}({u}_{f}[k])\in\mathcal{W}_{\ell 1},\,\forall\,\ell\in\mathcal{A}_{ui}\Longrightarrow u_{fi}[k]\in{\mathcal{U}}_{fi},\vspace{-3mm}$ (32)
for all $i\in\{1:N\}$. The following result employs the newly introduced sets in order to constrain the state variables of the first layer subcontrollers.
@@
Let the first layer subcontrollers be implemented as in (8)-(12). For each $i\in\{1:N\}$, define ${Z}_{ri}:=\mathrm{diag}(S_{ui},S_{xi})$ and the sets from (33) and (34a)-(34b), located at the top of the next page. Assume that:
@@ 2. $ii)$
Note that the $i^{\text{th}}$ area’s sets from (34a)-(34b) are computed efficiently, using only for those indices belonging to $\mathcal{N}_{i}$. This is owed to the sparsity structure displayed in (12), which serves to incentivise the synthesis of sparse control laws for the first layer (see [10]).
The result formalized via Proposition 5.1 will prove instrumental in guaranteeing the recursive feasibility of the second layer and, in retrospect, the choice to impose constraints upon $u_{fi}[k]$, instead of the $u_{i}[k]$ vectors in (26), becomes justified. Not only do the specialized implementations presented in (8) allow us to construct set-based descriptions (34a)-(34b) for the states belonging to the first layer implementations, but they also enable us to explicitly link the aforementioned sets to the second layer’s design objectives. By eliminating the need to individually select each of these sets (which is an intricate and highly empirical problem, in the distributed setting), the design procedure becomes greatly simplified.
@@ 5.4 Addressing Uncertainty in the Initial Conditions
Having obtained the means by which we will satisfy the inclusions given in (26), we now turn our attention to the disturbance discussed in Sections 4.2 and 4.3. In order to account for the inaccuracies modelled in (20), we begin by exploiting the expressions from (25) to compute
$\left\{\begin{split}\mathcal{H}_{xit}\subseteq\mathbb{R}^{n_{xi}}\text{ s.t. }\eta_{xi}[k_{0}+t]\in\mathcal{H}_{xit},\,\forall\,\tiny\begin{bmatrix}\nu_{x}^{\top}\ \nu_{w}^{\top}\end{bmatrix}^{\top}\in\mathcal{V},\\ \mathcal{H}_{uit}\subseteq\mathbb{R}^{n_{ui}}\text{ s.t. }\eta_{ui}[k_{0}+t]\in\mathcal{H}_{uit},\,\forall\,\tiny\begin{bmatrix}\nu_{x}^{\top}\ \nu_{w}^{\top}\end{bmatrix}^{\top}\in\mathcal{V},\end{split}\right.$ (35)
@@
$\widetilde{\mathcal{X}}_{it}:={\mathcal{X}}_{i}\ominus(-\mathcal{H}_{xit}),\vspace{-4mm}$ (36)
@@
$\widetilde{\mathcal{W}}_{\ell 1t}:=\mathcal{W}_{\ell 1}\ominus\big{(}-e^{\top}_{(\ell-\alpha_{ui})}\mathcal{H}_{uit}\big{)},\,\forall\,\ell\in\mathcal{A}_{ui}.\vspace{-3mm}$ (37)
Thus, our problem reduces to employing the dynamics from (21a)-(21b) with the aim of constraining $\widetilde{x}_{i}[k]$ and $\widetilde{u}_{fi}[k]$ to the sets obtained in (36) and (37), respectively. Recalling the identities from (24), this directly yields the desired inclusions from (26). All that remains is to compute bounding sets for the signals from (22)-(23).
@@ 5.5 Placing Constraints on Area Cross-coupling
$\mathrm{row}_{j}(w_{r\ell}[k])\in\mathcal{W}_{\ell j},\,\forall\,j\in\{1:n_{r\ell}\},\,\ell\in\mathcal{A}_{ui}\Longrightarrow\\ \Longrightarrow w_{i}[k]\in\mathcal{W}_{i},$ (38)
for all $i\in\{1:N\}$, by simply taking the Cartesian product of the $\mathcal{W}_{\ell j}$ sets appearing in (38). Since the aim of the second layer is to constrain the state variables of the network and those of the implementations given in (8) to the sets $\mathcal{X}_{i}$ and $\mathcal{W}_{i}$, respectively, we consider that the initial conditions of the first layer’s closed-loop system satisfy $x_{ci}\in\mathcal{X}_{i}$ along with $w_{ci}\in\mathcal{W}_{i}$, for all $i\in\{1:N\}$. In this case, (20) yields
$\widetilde{x}_{ci}={x}_{ci}+S^{\top}_{{xi}}\nu_{x}\in\widetilde{\mathcal{X}}_{ci}:=\mathcal{X}_{i}\oplus\tiny\begin{bmatrix}S^{\top}_{{xi}}&O\end{bmatrix}\mathcal{V},$ (39a)
$\widetilde{w}_{ci}={w}_{ci}+S^{\top}_{{wi}}\nu_{w}\in\widetilde{\mathcal{W}}_{ci}:=\mathcal{W}_{i}\oplus\tiny\begin{bmatrix}O&S^{\top}_{{wi}}\end{bmatrix}\mathcal{V},$ (39b)
@@
$\left\{\begin{split}&\widetilde{\Theta}_{xit}\subseteq\mathbb{R}^{n_{xi}}\text{ s.t. }\widetilde{\theta}_{xi}[k_{0}+t]\in\widetilde{\Theta}_{xit},\\ &\qquad\qquad\qquad\qquad\quad\forall\,\widetilde{x}_{cj}\in\widetilde{\mathcal{X}}_{cj},\ \widetilde{w}_{cj}\in\widetilde{\mathcal{W}}_{cj},\ j\in\mathcal{N}_{i},\\ &\widetilde{\Theta}_{uit}\subseteq\mathbb{R}^{n_{ui}}\text{ s.t. }\widetilde{\theta}_{ui}[k_{0}+t]\in\widetilde{\Theta}_{uit},\\ &\qquad\qquad\qquad\qquad\quad\forall\,\widetilde{x}_{cj}\in\widetilde{\mathcal{X}}_{cj},\ \widetilde{w}_{cj}\in\widetilde{\mathcal{W}}_{cj},\ j\in\mathcal{N}_{i},\end{split}\right.\normalsize\vspace{-2mm}$ (40)
@@
$\left\{\begin{split}&\widetilde{\Delta}_{xit}\subseteq\mathbb{R}^{n_{xi}}\text{ s.t. }\widetilde{\delta}_{xi}[k_{0}+t]\in\widetilde{\Delta}_{xit},\\ &\qquad\forall\,u_{s1j}[k]\in{\mathcal{U}}_{s1j},\,u_{s2j}[k]\in{\mathcal{U}}_{s2j},\,j\in\{1:N\}\backslash\{i\},\\ &\qquad\forall\,\widetilde{x}_{c\ell}\in\widetilde{\mathcal{X}}_{c\ell},\ \widetilde{w}_{c\ell}\in\widetilde{\mathcal{W}}_{c\ell},\ j\in\{1:N\}\backslash\mathcal{N}_{i},\\ &\widetilde{\Delta}_{uit}\subseteq\mathbb{R}^{n_{ui}}\text{ s.t. }\widetilde{\delta}_{ui}[k_{0}+t]\in\widetilde{\Delta}_{uit},\\ &\qquad\forall\,u_{s1j}[k]\in{\mathcal{U}}_{s1j},\,u_{s2j}[k]\in{\mathcal{U}}_{s2j},\,j\in\{1:N\}\backslash\{i\},\\ &\qquad\forall\,\widetilde{x}_{c\ell}\in\widetilde{\mathcal{X}}_{c\ell},\ \widetilde{w}_{c\ell}\in\widetilde{\mathcal{W}}_{c\ell},\ \ell\in\{1:N\}\backslash\mathcal{N}_{i}.\end{split}\right.\normalsize\vspace{-3mm}$ (41)
@@
By inspecting Figure 1 along with the various sets defined in this section, the intended use of $u_{s1i}[k]$ and $u_{s2i}[k]$ along with the trade-off associated with selecting the sets from (30a)-(30b) become clear. Just as $u_{s2i}[k]$ is primarily meant to steer the $i^{\text{th}}$ area’s state into the desired set, so too is $u_{s1i}[k]$ chiefly responsible for keeping the state variables of the respective area’s first layer subcontroller (and, implicitly, its command signal vector) within appropriate bounds. Thus, in choosing the sets from (30b), one negotiates a command-budgeting balance (as limited by $\mathcal{U}_{i}$) between the first and second layers. On the other hand, the more freedom to act $u_{s1i}[k]$ is given, via the set from (30a), its effect will expand throughout the sets from (33) and (34a)-(34a). Similarly, if the network’s areas are not all perfectly decoupled (recall Remark 4.1), then too much leeway given to $u_{s1i}[k]$ and $u_{s2i}[k]$ may result in significant area cross-coupling.
$\scriptsize\begin{split}&\min\limits_{u_{s1i}[k_{0}],\dots,u_{s1i}[k_{0}+T_{i}+\overline{T}_{i}-1],u_{s2i}[k_{0}],\dots,u_{s2i}[k_{0}+T_{i}+\overline{T}_{i}-1]}\sum_{t=1}^{T_{i}+\overline{T}_{i}}\widetilde{q}_{i}^{\top}[t]Q_{i}[t]\widetilde{q}_{i}[t]+c_{i}[t]^{\top}\widetilde{q}_{i}[t]+r_{i}[t],\\ &\text{subject to }\left\{\begin{split}&\widetilde{q}_{i}[t]=\scriptsize\begin{bmatrix}\xi_{i}^{\top}[k_{0}+t]\ \ u_{s1i}^{\top}[k_{0}+t-1]\ \ u_{s2i}^{\top}[k_{0}+t-1]\end{bmatrix}^{\top},\,\forall\,t\in\{1:T_{i}+\overline{T}_{i}\},\\ &\xi_{i}[k_{0}+t]=A_{si}\xi_{i}[k_{0}+t-1]+B_{s1i}u_{s1i}[k_{0}+t-1]+B_{s2i}u_{s2i}[k_{0}+t-1],\,\forall\,t\in\{1:T_{i}+\overline{T}_{i}\},\\ &\xi_{i}[k_{0}+t]\in\Xi_{it},\,\forall\,t\in\{1:T_{i}\},\ \xi_{i}[k_{0}]=0,\\ &u_{s1i}[k_{0}+t-1]\in{\mathcal{U}}_{s1i},\,u_{s2i}[k_{0}+t-1]\in{\mathcal{U}}_{s2i},\,\forall\,t\in\{1:T_{i}\}.\end{split}\right.\end{split}\normalsize$ (42)
$\Xi_{it}:=\left\{\phi\in\mathbb{R}^{n_{si}}\ \left|{\ \scriptsize\begin{bmatrix}H_{1it}C_{xi}\\ H_{2\left(\alpha_{ui}+1\right)t}\mathrm{row}_{1}(C_{ui})\\ \vdots\\ H_{2\left(\alpha_{ui}+n_{ui}\right)t}\mathrm{row}_{n_{ui}}(C_{ui})\end{bmatrix}}\right.\phi\leq\scriptsize\begin{bmatrix}h_{1it}-H_{1it}\widetilde{\theta}_{xi}[k_{0}+t]\\ h_{2\left(\alpha_{ui}+1\right)t}-H_{2\left(\alpha_{ui}+1\right)t}\mathrm{row}_{1}\left(\widetilde{\theta}_{ui}[k_{0}+t]\right)\\ \vdots\\ h_{2\left(\alpha_{ui}+n_{ui}\right)t}-H_{2\left(\alpha_{ui}+n_{ui}\right)t}\mathrm{row}_{n_{ui}}\left(\widetilde{\theta}_{ui}[k_{0}+t]\right)\end{bmatrix}\right\},\forall\,t\in\{1:T_{i}\}.\vspace{-1mm}$ (43)
With all of these sets now defined according to their corresponding areas, objective II) from Section 3.3 has been successfully addressed. Therefore, we now move on to the following section, in which we state the formulations of our distributed MPC controllers that ensure the (recursive) satisfaction of the second layer’s objectives.

### v4

@@ 5.1 Placing Constraints on Exogenous Disturbance
$\left\{\hskip-2.84526pt\scriptsize\begin{array}[]{ll}\mathcal{D}_{d}\subseteq\mathbb{R}^{n_{d}}\,\text{s.t. }d[k]\in\mathcal{D}_{d},&\mathcal{D}_{\beta_{f}}\subseteq\mathbb{R}^{n_{u}}\,\text{s.t. }\beta_{f}[k]\in\mathcal{D}_{\beta_{f}},\\ \mathcal{D}_{\beta_{s1}}\subseteq\mathbb{R}^{n_{x}}\,\text{s.t. }\beta_{s1}[k]\in\mathcal{D}_{\beta_{s1}},&\mathcal{D}_{\beta_{s2}}\subseteq\mathbb{R}^{n_{x}}\,\text{s.t. }\beta_{s2}[k]\in\mathcal{D}_{\beta_{s2}}.\end{array}\right$. (27)
@@
[…]
Remark 5.2.
One of the key benefits of our two-layer approach is the fact that, from the point of view of the NRF-based layer, the signals $u_{s1}[k]$ and $u_{s2}[k]$ are perceived in the same manner as the all the other components the vector from (7), namely as additive and exogenous disturbance. Consequently, the special disturbance attenuation properties (such as string stability, see [plutonizare] and [SS_rev], along with the references therein) of our control system are preserved when adding the second layer on top of the first. Note that such special properties are particularly challenging to ensure using exclusively MPC-based techniques, a topic which has generated a great deal of research in system-theoretical literature [plat_MPC_1, plat_MPC_2, plat_MPC_3, plat_MPC_4, plat_MPC_5]. One of the main contributions of the present work is the fact that it combines the native benefits of two separate techniques, disturbance attenuation for NRF and constraint management for MPC, into a single framework.
Moving onward, we address the variables associated with the first-layer subcontrollers and their implementations.
@@ 5.2 Placing Constraints on First Layer Commands
$u_{i}[k]=u_{fi}[k]+u_{s2i}[k]+S^{\top}_{{ui}}\beta_{s2}[k],\,\forall\,i\in\{1:N\}.\vskip-8.53581pt$ (29)
For the state dynamics obtained in Section 4, the problem of constraining the command signals sent to the network reduces to judiciously constraining the control effort of both the first and the second layer. Thus, for each area, we choose the pair of sets introduced in (8), namely
@@
${\mathcal{U}}_{fi}:=\mathcal{U}_{i}\ominus(\,\mathcal{U}_{s2i}\oplus(S^{\top}_{{ui}}\mathcal{D}_{\beta_{s2}})),\,\forall\,i\in\{1:N\}.\vskip-8.53581pt$ (31)
@@
The freely chosen sets given in (30a)-(30b) are the main design parameters of the second layer. They can be assimilated with the extent to which the second layer is permitted to interfere in the dynamics of the first layer’s closed-loop system, so as to ensure that (26) is satisfied. The manner in which these sets may be chosen, and the resulting trade-off, will be addressed in the sequel.
5.3 Placing Constraints on First-Layer States
@@
Remark 5.4.
The results presented in this section represent one of the key theoretical novelties introduced in this paper. In the classical tube MPC setting (see, for example, [R8A, R8B, R8D]), the command constraints on the signals analogous to $u_{fi}[k]$ are straightforward to compute by means of constant matrices. In our framework, however, it becomes necessary to pass through the dynamics in (10), which significantly complicates the resulting expressions. Given the delicate nature of this undertaking, said expressions are presented, within the current section, in as explicit a manner as possible, regardless of the apparent similarity with standard tube MPC (recall the first paragraph of Section 4). Nevertheless, the benefits of NRF-based control (as stated in Remark 5.2) more than make up for the increase in complexity when formulating the synthesis procedure (and not, we point out, when implementing the control laws). Indeed, a further advantage of our approach is the fact that, instead of limiting ourselves to block-diagonal and constant stabilising feedback (see Assumption 1 in [R8A, R8B, R8D]), we leverage a greater degree of freedom in considering dynamical stabilising feedback (10) endowed with generic sparsity patterns (11).
We begin by inspecting the distinctive structure of the matrices which make up the realisations from (10). The latter enables us to naturally partition each of the state vectors belonging to the first-layer subcontrollers into $w_{i}[k]=\tiny\begin{bmatrix}w_{r(\alpha_{ui}+1)}^{\top}[k]&\dots&w_{r(\alpha_{ui}+n_{ui})}^{\top}[k]\end{bmatrix}$, in which the vectors $w_{r\ell}[k]$ have $n_{r\ell}$ rows for all $\ell\in\{1:n_{u}\}$, i.e., the same as the matrices $A_{r\ell}$ and $B_{r\ell}$ from (11). Recalling (11), the key observation which underpins the entire mathematical mechanism presented in this section is the fact that $\mathrm{row}_{\ell}(u_{f}[k])=C_{r\ell}w_{r\ell}[k]=\mathrm{row}_{1}(w_{r\ell}[k])$, for all $\ell\in\{1:n_{u}\}$. Thus, our current objective becomes apparent: we aim to employ the bounds placed upon the partitions of $u_{f}[k]$ to satisfy further set inclusions with respect to $w_{r\ell}[k]$, in a distributed manner.
@@
$\mathrm{row}_{\ell}({u}_{f}[k])\in\mathcal{W}_{\ell 1},\,\forall\,\ell\in\mathcal{A}_{ui}\Longrightarrow u_{fi}[k]\in{\mathcal{U}}_{fi},\vskip-8.53581pt$ (32)
for all $i\in\{1:N\}$. The following result employs the newly introduced sets with the aim of constraining the state variables of the first-layer subcontrollers.
@@
Let the first-layer subcontrollers be implemented as in (9)-(13). For each $i\in\{1:N\}$, define ${Z}_{ri}:=\mathrm{diag}(S_{ui},S_{xi})$ and the sets from (31) and (34a)-(34b), located at the top of this page. Assume that:
@@ 3. A3)
Note that the $i^{\text{th}}$ area’s sets given in (34a)-(34b) are computed efficiently, using only for those indices belonging to $\mathcal{N}_{i}$. This is owed to the sparsity structure displayed in (13), which serves to incentivise the synthesis of sparse control laws for the first layer (see Appendix A, along with [part1]). Note that the bounds associated with these sets could also be tightened by adapting the ideas presented in [R8B], through the sequential reduction of those $\mathcal{X}_{j}$, $\mathcal{U}_{j}$, $\mathcal{U}_{s1j}$ and $\mathcal{U}_{s2j}$ for which $j\in\mathcal{N}_{i}$.
The result formalised via Proposition 5.1 will prove instrumental in guaranteeing the recursive feasibility of the second layer and, in retrospect, the choice to impose constraints upon $u_{fi}[k]$, instead of the $u_{i}[k]$ vectors in (26), becomes justified. Not only do the specialised implementations presented in (10) allow us to construct set-based descriptions (34a)-(34b) for the states belonging to the first-layer implementations, but they also enable us to explicitly link the sets mentioned above to the second layer’s design objectives. By eliminating the need to individually select each of these sets (which is an intricate and highly empirical problem, in the distributed setting), the design procedure becomes greatly simplified.
Remark 5.6.
By inspecting Figure 1 along with the various sets defined in this section, the intended use of $u_{s1i}[k]$ and $u_{s2i}[k]$ becomes clear. Indeed, just as $u_{s2i}[k]$ is primarily meant to ensure that $x_{i}[k]\in\mathcal{X}_{i}$, so too is $u_{s1i}[k]$ chiefly responsible for ensuring that $\mathrm{row}_{j}(w_{r\ell}[k])\in\mathcal{W}_{\ell j}$ and, implicitly, that $u_{i}[k]\in\mathcal{U}_{i}$ (recall Proposition 5.1).
@@ 5.4 Addressing Uncertainty in the Initial Conditions
Having obtained the means by which we will satisfy the inclusions given in (26), we now turn our attention to the disturbance discussed in Sections 4.3 and 4.4. In order to account for the inaccuracies modelled in (20), we begin by exploiting the expressions from (25) to compute
[…]
@@
$\widetilde{\mathcal{X}}_{it}:={\mathcal{X}}_{i}\ominus(-\mathcal{H}_{xit}),\vskip-14.22636pt$ (36)
@@
$\widetilde{\mathcal{W}}_{\ell 1t}:=\mathcal{W}_{\ell 1}\ominus\big(-e^{\top}_{(\ell-\alpha_{ui})}\mathcal{H}_{uit}\big),\,\forall\,\ell\in\mathcal{A}_{ui}.\vskip-8.53581pt$ (37)
Thus, our problem reduces to employing the dynamics from (21a)-(21b) with the aim of constraining $\widetilde{x}_{i}[k]$ and $\widetilde{u}_{fi}[k]$ to the sets obtained in (36) and (37), respectively. Recalling the identities from (24) directly yields the desired inclusions from (26). All that remains is to compute bounding sets for the signals from (22)-(23).
$\Xi_{it}:=\left\{\phi\in\mathbb{R}^{n_{si}}\ \left|{\ \scriptsize\begin{bmatrix}H_{1it}C_{xi}\\ H_{2\left(\alpha_{ui}+1\right)t}\mathrm{row}_{1}(C_{ui})\\ \vdots\\ H_{2\left(\alpha_{ui}+n_{ui}\right)t}\mathrm{row}_{n_{ui}}(C_{ui})\end{bmatrix}}\right.\phi\leq\scriptsize\begin{bmatrix}h_{1it}-H_{1it}\widetilde{\theta}_{xi}[k_{0}+t]\\ h_{2\left(\alpha_{ui}+1\right)t}-H_{2\left(\alpha_{ui}+1\right)t}\mathrm{row}_{1}\left(\widetilde{\theta}_{ui}[k_{0}+t]\right)\\ \vdots\\ h_{2\left(\alpha_{ui}+n_{ui}\right)t}-H_{2\left(\alpha_{ui}+n_{ui}\right)t}\mathrm{row}_{n_{ui}}\left(\widetilde{\theta}_{ui}[k_{0}+t]\right)\end{bmatrix}\right\},\forall\,t\in\{1:T_{i}\}.\vskip-2.84526pt$ (44)
@@ 5.5 Placing Constraints on Area Cross-coupling
$\hskip-5.69054pt\scriptsize\begin{array}[]{l}\mathrm{row}_{j}(w_{r\ell}[k])\in\mathcal{W}_{\ell j},\forall j\in\{1:n_{r\ell}\},\ell\in\mathcal{A}_{ui}\Rightarrow w_{i}[k]\in\mathcal{W}_{i},\end{array}\normalsize\hskip-2.84526pt\vskip-8.53581pt$ (38)
for all $i\in\{1:N\}$, by simply taking the Cartesian product of the $\mathcal{W}_{\ell j}$ sets appearing in (38). Since the aim of the second layer is to constrain the state variables of the network and those of the implementations given in (10) to the sets $\mathcal{X}_{i}$ and $\mathcal{W}_{i}$, respectively, we consider that the initial conditions of the first layer’s closed-loop system satisfy $x_{ci}\in\mathcal{X}_{i}$ along with $w_{ci}\in\mathcal{W}_{i}$, for all $i\in\{1:N\}$. In this case, the definitions introduced in (20) yield
$\hskip-2.84526pt\widetilde{x}_{ci}={x}_{ci}+S^{\top}_{{xi}}\nu_{x}\in\widetilde{\mathcal{X}}_{ci}:=\mathcal{X}_{i}\oplus\tiny\begin{bmatrix}S^{\top}_{{xi}}&O\end{bmatrix}\mathcal{V},\hskip-2.84526pt$ (39a)
$\hskip-2.84526pt\widetilde{w}_{ci}={w}_{ci}+S^{\top}_{{wi}}\nu_{w}\in\widetilde{\mathcal{W}}_{ci}:=\mathcal{W}_{i}\oplus\tiny\begin{bmatrix}O&S^{\top}_{{wi}}\end{bmatrix}\mathcal{V},\hskip-2.84526pt$ (39b)
@@
[…]
@@
[…]
@@
Note that, in choosing the sets from (30b), one negotiates a command-budgeting balance (as limited by $\mathcal{U}_{i}$) between the first and second layers. This induces a trade-off associated with selecting the sets from (30a)-(30b), since the more freedom to act $u_{s1i}[k]$ is given, via the set from (30a), its effect will expand throughout the sets from (31) and (34a)-(34b). Similarly, if the network’s areas are not all perfectly decoupled (recall Remark 4.4), then too much leeway given to $u_{s1i}[k]$ and $u_{s2i}[k]$ may result in significant area cross-coupling. In this case, one may derive an extended version of the technique from [R8C] (which manifests its procedure’s effectiveness on only two collaborating areas) to mitigate these particular aspects. Section 4 in [R8C] also provides valuable insight into negotiating this particular type of trade-off.
With all of these sets now defined according to their corresponding areas, we now move on to the following section, in which we present guarantees for the (recursive) satisfaction of the second layer’s constraints.